# From the Stochastic Embedding Sufficiency Theorem to a Superspace Diffusion Framework

### 20 Mar 2026 v1

@@ Abstract
A blind recovery protocol, receiving only raw time series and sampling interval, is applied identically to nine physical domains: classical mechanics, statistical mechanics, nuclear physics, quantum mechanics, chemical kinetics, electromagnetism, relativistic quantum mechanics, quantum harmonic oscillator dynamics, and quantum electrodynamics. The pipeline recovers the governing equations of each domain with errors ranging from 0.026% to approximately 1%, with no null hypothesis rejected at the 5% significance level. Fundamental physical constants emerge in both the drift and diffusion channels without prior specification.
The recovered diffusion coefficients, viewed across domains, constitute an empirical pattern, the $\sigma$-continuum, in which the Planck constant, the Boltzmann constant, and the speed of light each play structurally distinct roles. Interpretation of the relativistic and quantum field theory results, combined with three independent uniqueness arguments, determines the gravitational diffusion coefficient as one Planck length per square root of Planck time, non-parametrically derived from first principles.
A set of four canonical axioms formalises the framework. Physical time emerges as a monotone functional of the stochastic evolution. Within these axioms and the short-memory limit, the drift, covariance operator, and fluctuation amplitude are all fixed. The resulting superspace diffusion hypothesis contains non-parametric, first-principles predictions within the stated axioms and generates falsifiable predictions against galactic kinematic data as developed in a companion paper (Part II).
@@ 1.3 Scope and Structure
Interpretation of the relativistic and quantum field theory results yields a specific superspace diffusion hypothesis, formalised by four canonical axioms within which the gravitational diffusion coefficient, the drift, the covariance operator, and physical time itself are all fixed. The resulting framework contains non-parametric, first-principles predictions within the stated axioms. These predictions are tested against observational galactic kinematic data (rotation curves, velocity dispersions, wide binaries, galaxy clusters, gravitational lensing, and cosmic voids) in a companion paper (Part II).
The paper is structured as follows. Section 2 describes the recovery pipeline, the blind protocol, the data generation procedures, and the statistical validation framework. Section 3 presents the nine-domain results with statistical rigour and introduces the $\sigma$-continuum. Section 4 interprets the results in three stages: relativistic structure, the gravitational diffusion coefficient, and axiomatic formalisation. Section 5 states the conclusions. Appendices A–D provide the mathematical derivations: the Stochastic Embedding Sufficiency Theorem (A), the relativistic transformation properties (B), the gravitational diffusion coefficient (C), and the uniqueness theorems (D).
@@ 2.2 The Stochastic Embedding Pipeline
where $d$ is the embedding dimension and $\tau$ is the delay. The embedding dimension is selected using the $E_{1}$ statistic of Cao [15]: $E_{1}(d)$ measures the ratio of nearest-neighbour distances in successive embedding dimensions, and saturation of $E_{1}$ indicates that the correlation manifold has been unfolded. In all experiments reported here, $\tau=1$ (one sampling interval) is used, where the embedding dimension $d$ is domain-specific and determined by $E_{1}$ saturation.
@@ 2.4 Data Generation
Table 1: Data generation summary for the nine domain validations. NN denotes the number of time series samples provided to the pipeline; $k$ denotes the number of nearest neighbours used in Stage 2. All data is synthetically generated from known physics. Theoretical diffusion coefficients are determined by the governing equations of each domain.
Test Domain Simulation Theoretical $\sigma$ NN $k$
C1 Classical mechanics Numerical integration of Kepler orbit with GR precession 0 (deterministic) 36 525 20
C2 Statistical mechanics Euler–Maruyama Langevin equation $\sqrt{2\gamma k_{B}T/m}$ 100 000 200
C3 Nuclear physics Poisson sampling $\sqrt{\mu}$ (Fano =1=1) 10 000 200
C4 Quantum mechanics Spectral Schrödinger evolution (wavepacket width) $\sqrt{\hbar/m}$ 2 001 30
C5 Chemical kinetics Tau-leaping birth–death process $\propto\Omega^{-1/2}$ 1 001 30
C6 Electromagnetism Finite-difference Maxwell wave equation 0 (deterministic) 30
C7 Relativistic QM Euler–Maruyama Nelson SDE with relativistic dispersion $\sigma_{0}/\gamma$ 2 000 30
C8 Quantum field theory Euler–Maruyama damped QHO at zero-point $\sqrt{\hbar\omega/2}$ 100
C9 QED (photon field) Euler–Maruyama QHO with $m_{\mathrm{eff}}=\hbar\omega/c^{2}$ $c\sqrt{\gamma}$ 100
@@ 2.5 Statistical Validation Framework
Hypothesis testing assesses whether the recovered value is statistically consistent with the theoretical prediction. For each domain, the null hypothesis $H_{0}\!:\hat{\theta}=\theta_{\mathrm{true}}$ is tested using the appropriate test statistic (Student’s $t$ or equivalent); the resulting $p$-value is reported. A $p$-value exceeding $0.05$ indicates that the null hypothesis is not rejected at the 5% significance level.
Confidence intervals (95%) are constructed from 50 independent pipeline runs, each with independently generated synthetic data. Bootstrap resampling (2 000 resamples) is used to verify that analytical standard errors are consistent with empirical uncertainty [20].
@@
Where multiple tests are combined, the Bonferroni correction is applied to control the family-wise error rate.
@@ 3.1 Deterministic Domains: $\sigma=0$
For the cumulative precession series (Fig. 2a,b), SVD identified rank 1, and the recovered drift reproduced the theoretical precession rate to 0.02%. The median recovered diffusion was $\hat{\sigma}=7.19\times 10^{-14}$ arcsec, indistinguishable from zero at machine precision.
@@
The Planck constant was extracted from the recovered drift: $\hat{\hbar}=1.00026$ (error $0.026\%$ relative to the simulation value $\hbar=1$). Across six independent simulations with initial widths $\sigma_{0}\in[1,5]$, the recovered $\hat{\hbar}$ was stable to within $0.05\%$ (Fig. 3d). Fifty independent pipeline runs with added measurement noise yielded median $\hat{\hbar}=1.00053$ with 95% confidence interval $[0.997,1.004]$ (Fig. 3e).
@@
A classical electromagnetic wave was propagated using a finite-difference scheme. The pipeline recovered $\hat{\sigma}=3.51\times 10^{-14}$ (indistinguishable from zero) and a propagation speed $\hat{c}/c_{\mathrm{true}}=1.0000$ (Fig. 8a,b). Electromagnetism, like classical mechanics, operates in the deterministic limit $\sigma=0$. The C6 figure panels are presented together with C7 in Fig. 8.
@@ 3.2 Additive Noise: Constant $\sigma$
The pipeline recovered a linear drift $\hat{\mu}(v)=-\hat{\gamma}v$ (Fig. 4c) and a flat diffusion profile $\hat{\sigma}(v)\approx\mathrm{const.}$ (Fig. 4b), consistent with additive noise. The diffusion coefficient was recovered to 0.79% error. The Boltzmann constant, extracted from the fluctuation–dissipation relation $k_{B}=m\hat{\sigma}^{2}/(2\hat{\gamma}T)$ [5, 22], was recovered as $\hat{k}_{B}/k_{B}=0.984$ with 95% CI [0.955, 1.016][0.955,\,1.016] from 50 independent simulations (Fig. 4f). The null hypothesis $\hat{k}_{B}=k_{B}$ was not rejected.
@@ 3.3 State-Dependent Diffusion
The pipeline recovered a state-dependent diffusion profile $\hat{\sigma}^{2}(\mu)$ lying on the Poisson line $\sigma^{2}=\mu$ across a two-decade range of mean count levels (Fig. 5b). On a log–log plot, the variance–mean slope was $1.0002$ (theoretical value $1.0000$; error $0.02\%$; $R^{2}=0.9997$). Fifty independent simulations at $\mu=3000$ yielded a Fano factor of median $1.0000$ with 95% CI [0.961, 1.030][0.961,\,1.030] and $p=0.50$ for the null hypothesis $F=1$ (Fig. 5e). The 250-realisation ensemble achieved 95.8% CI coverage.
@@
The pipeline recovered $\hat{\sigma}$ to 0.35% error (Fig. 6b). Across nine frequency values spanning a decade, the recovered $\hat{\sigma}^{2}$ was proportional to $\omega$ with $R^{2}=0.9999$ (Fig. 6c), confirming the linear frequency dependence predicted by the fluctuation–dissipation relation. Twenty independent random seeds at the primary frequency yielded errors uniformly below 1% (Fig. 6d).
@@ 3.4 State-Dependent Drift and Diffusion
The pipeline was applied independently at seven system sizes ($\Omega\in[50,5000]$). On a log–log plot of recovered $\hat{\sigma}$ versus $\Omega$ (Fig. 7d), linear regression yielded a scaling exponent $\hat{\alpha}=-0.502\pm 0.001$ (theoretical $\alpha=-0.500$; error $0.38\%$; $R^{2}=1.0000$). The null hypothesis $\hat{\alpha}=-0.500$ was not rejected ($p=0.46$). Bootstrap resampling from 50 independent pipeline runs gave $\hat{\alpha}=-0.501\pm 0.006$ (Fig. 7e). The 250-realisation ensemble at $\Omega=500$ achieved 94.0% CI coverage.
@@ 3.5 Relativistic Corrections
For C6, a classical electromagnetic wavepacket was propagated by finite-difference integration. The pipeline recovered $\hat{\sigma}=3.51\times 10^{-14}$ (indistinguishable from zero) and a propagation speed $\hat{c}/c_{\mathrm{true}}=1.0000$ (Fig. 8a,b).
@@
On a log–log plot of recovered $\hat{\sigma}/\hat{\sigma}_{0}$ versus $\gamma$ (Fig. 8d), the slope was 0.993-0.993 with 95% CI [1.018,0.970][-1.018,-0.970] (theoretical 1.000-1.000; error $0.66\%$; $R^{2}=0.9999$). The null hypothesis of unit slope was not rejected ($p=0.58$).
@@
The two factors operate at different levels (noise amplitude versus drift), arise from different inputs (special relativity versus Klein–Gordon current conservation), and use different mathematical tools (DDS theorem versus Fokker–Planck self-consistency). Their product $\sigma_{\mathrm{KG}}=\sigma_{0}/\gamma$ is confirmed by the C7 recovery: slope 0.993-0.993 with 95% CI [1.018,0.970][-1.018,-0.970], consistent with the predicted exponent 1-1 ($p=0.58$). The independent recovery of the group velocity $v_{g}$ at each momentum mitigates the potential circularity that $\gamma$ depends on $v$: both channels: drift and diffusion: carry relativistic content, recovered independently from the same blind pipeline.

### 25 Mar 2026 v2

@@ Abstract
A blind recovery protocol, receiving only raw time series and sampling interval, is applied identically to nine physical domains: classical mechanics, statistical mechanics, nuclear physics, quantum mechanics, chemical kinetics, electromagnetism, relativistic quantum mechanics, quantum harmonic oscillator dynamics, and quantum electrodynamics. The pipeline recovers the governing equations of each domain blindly, with recovery errors quantified by bootstrap confidence intervals from 50 independent pipeline runs per domain. Fundamental physical constants (the Boltzmann constant, the Planck constant, the speed of light, the Fano factor, and the Van Kampen scaling exponent) emerge in both the drift and diffusion channels without prior specification.
The recovered diffusion coefficients, viewed across domains, constitute an empirical pattern, the $\sigma$-continuum, in which $k_{B}$, $\hbar$, and $c$ each play structurally distinct roles. The Gravitational Diffusion Theorem, derived from the fluctuation–dissipation theorem, the massless mode structure of linearised gravity, and gravitational self-coupling via the equivalence principle, each among the most precisely tested results in physics, determines the gravitational diffusion coefficient as one Planck length per square root of Planck time, with three independent uniqueness arguments converging on the dimensionless prefactor $\alpha=1$.
Four canonical axioms formalise the framework: three grounded in standard physics (Wheeler’s superspace, the derived fluctuation amplitude, and classical correspondence with general relativity) and one interpretive choice (epistemic probability). Physical time emerges as a monotone functional of the stochastic evolution. Within these axioms, the noise character, drift, covariance operator, and fluctuation amplitude are all uniquely determined by theorem (Appendix D), yielding the superspace diffusion hypothesis:
$\mathrm{d}g_{ij}=\mathcal{D}_{ij}[g]\,\mathrm{d}\tau+\ell_{P}\,\mathrm{d}W_{ij},$
where all coefficients are non-parametric, first-principles consequences of the axioms. Coarse-graining of the superspace Fokker–Planck equation via the Mori–Zwanzig projection yields predictions for gravitational acceleration at galactic scales that are testable against kinematic data.
@@ 1.3 Scope and Structure
Derivation from the relativistic and quantum field theory results yields a specific superspace diffusion hypothesis, formalised by four canonical axioms within which the gravitational diffusion coefficient, the drift, the covariance operator, and physical time itself are all fixed. The resulting framework contains non-parametric, first-principles predictions within the stated axioms. At galactic scales, the Fokker–Planck equation on superspace generates predictions for gravitational acceleration that are testable against kinematic data including rotation curves, velocity dispersions, wide binaries, galaxy clusters, gravitational lensing, and cosmic voids.
The paper is structured as follows. Section 2 describes the recovery pipeline, the blind protocol, the data generation procedures, and the statistical validation framework. Section 3 presents the nine-domain results with statistical rigour and introduces the $\sigma$-continuum. Section 4 derives the gravitational diffusion coefficient from the results in three stages: relativistic structure (§4.2, summarising Appendix B), the gravitational diffusion coefficient (§4.3, summarising Appendix C), and axiomatic formalisation with uniqueness proofs (§4.4, summarising Appendix D). Section 5 states the conclusions. Appendices A–D provide the mathematical derivations: the Stochastic Embedding Sufficiency Theorem (A), the relativistic transformation properties (B), the gravitational diffusion coefficient (C), and the uniqueness theorems (D).
@@ 2.2 The Stochastic Embedding Pipeline
where $d$ is the embedding dimension and $\tau$ is the delay. The embedding dimension is selected using the $E_{1}$ statistic of Cao [15]: $E_{1}(d)$ measures the ratio of nearest-neighbour distances in successive embedding dimensions, and saturation of $E_{1}$ indicates that the correlation manifold has been unfolded. For non-oscillatory systems (C1–C7), $\tau=1$ (one sampling interval) is used. For oscillatory systems (C8, C9), where the sampling interval is much shorter than the dynamical timescale, $\tau$ is set to the first zero-crossing of the autocorrelation function, ensuring that the delay embedding spans a meaningful portion of the phase space (Table 1). The embedding dimension $d$ is domain-specific and determined by $E_{1}$ saturation.
@@ 2.4 Data Generation
Table 1: Data generation summary for the nine domain validations. NN denotes the number of time series samples provided to the pipeline; $k$ denotes the number of nearest neighbours used in Stage 2; $\tau_{\mathrm{embed}}$ is the embedding lag in samples. For oscillatory systems (C8, C9), $\tau_{\mathrm{embed}}$ is set to the first zero-crossing of the autocorrelation function rather than unity, ensuring that the delay-coordinate embedding spans a meaningful portion of the dynamical phase space. All data is synthetically generated from known physics.
Test Domain Simulation Theoretical $\sigma$ NN $k$ $\tau_{\mathrm{embed}}$
C1 Classical mechanics Numerical integration of Kepler orbit with GR precession 0 (deterministic) 36 525 20 1
C2 Statistical mechanics Euler–Maruyama Langevin equation $\sqrt{2\gamma k_{B}T/m}$ 100 000 200 1
C3 Nuclear physics Poisson sampling $\sqrt{\mu}$ (Fano =1=1) 10 000 200 1
C4 Quantum mechanics Spectral Schrödinger evolution (wavepacket width) $\sqrt{\hbar/m}$ 2 001 30 1
C5 Chemical kinetics Tau-leaping birth–death process $\propto\Omega^{-1/2}$ 1 001 30 1
C6 Electromagnetism Gaussian wavepacket propagation (centroid tracking) 0 (deterministic) 500 30 1
C7 Relativistic QM Euler–Maruyama Nelson SDE with relativistic dispersion $\sigma_{0}/\gamma$ 2 000 30 1
C8 Quantum field theory Euler–Maruyama damped QHO at zero-point $\sqrt{\gamma\hbar\omega/m}$ 20 000 100 $\tau_{\mathrm{ACF}}$
C9 QED (photon field) Euler–Maruyama QHO with $m_{\mathrm{eff}}=\hbar\omega/c^{2}$ $c\sqrt{\gamma}$ 20 000 100 $\tau_{\mathrm{ACF}}$
@@ 2.5 Statistical Validation Framework
Bootstrap confidence intervals (95%) are the primary statistical assessment. For each domain, 50 independent pipeline runs are performed, each with independently generated synthetic data. Each run yields a normalised ratio $\hat{\theta}/\theta_{\mathrm{true}}$. The pairwise bootstrap (10 000 resamples of the 50 ratios with replacement) gives a distribution of the sample median; the 2.5th and 97.5th percentiles of this distribution define the 95% confidence interval [20]. If the interval contains 1.0, the recovered quantity is consistent with theory at the stated confidence level.
For each domain, both the drift $\hat{\mu}$ and diffusion $\hat{\sigma}$ are extracted and reported where the physics admits a meaningful comparison (dual-channel reporting). For deterministic systems (C1, C4, C6), the drift channel recovers the governing dynamical parameter while the diffusion channel confirms $\hat{\sigma}/\mathrm{signal}\ll 1$. For stochastic systems, the diffusion channel recovers $\sigma$ or a derived physical constant, and the drift channel recovers the deterministic component of the dynamics.
For domains testing scaling exponents (C5: Van Kampen exponent; C7: relativistic suppression slope), the pipeline is run at multiple parameter values within each simulation, and the log–log slope is computed per simulation before bootstrapping. The multiplicative $k$-NN estimator bias, which is approximately uniform across the parameter range, cancels in the slope [14].
@@ 3.1 Deterministic Domains: $\sigma=0$
For the cumulative precession series (Fig. 2a,b), SVD identified rank 1, and the recovered drift from a single primary run reproduced the theoretical precession rate to 0.02%. The median recovered diffusion was $\hat{\sigma}=7.19\times 10^{-14}$ arcsec, indistinguishable from zero at machine precision. Across 50 independent pipeline runs with small measurement noise, the drift-channel ratio $\hat{\theta}/\theta_{\mathrm{true}}$ had median $1.0000$ (within 1 ppm of unity) with bootstrap 95% CI [1.0000, 1.0000][1.0000,\,1.0000]; the diffusion channel confirmed $\hat{\sigma}/\mathrm{signal}=1.2\times 10^{-6}$.
@@
The Planck constant was extracted from the recovered drift: $\hat{\hbar}=1.00026$ (error $0.026\%$ relative to the simulation value $\hbar=1$). Across six independent simulations with initial widths $\sigma_{0}\in[1,5]$, the recovered $\hat{\hbar}$ was stable to within $0.05\%$ (Fig. 3d). Fifty independent pipeline runs with added measurement noise yielded median $\hat{\hbar}/\hbar=1.0005$ with bootstrap 95% CI [1.0001, 1.0015][1.0001,\,1.0015] (Fig. 3e); the diffusion channel confirmed $\hat{\sigma}/\mathrm{signal}=9.5\times 10^{-3}$ (deterministic).
@@
The electromagnetic wave propagation results are presented jointly with C7 (Klein–Gordon suppression) in §3.5 below, as the transition from deterministic wave propagation ($\sigma=0$) to relativistic diffusion suppression constitutes the pivotal connection between classical and quantum regimes. In summary: the pipeline recovered $\hat{\sigma}/\mathrm{signal}=4.3\times 10^{-7}$ and $\hat{c}/c_{\mathrm{true}}=1.0000$ with bootstrap 95% CI [1.0000, 1.0000][1.0000,\,1.0000] from 50 independent runs (Fig. 8a,b).
@@ 3.2 Additive Noise: Constant $\sigma$
The pipeline recovered a linear drift $\hat{\mu}(v)=-\hat{\gamma}v$ (Fig. 4c) and a flat diffusion profile $\hat{\sigma}(v)\approx\mathrm{const.}$ (Fig. 4b), consistent with additive noise. The Boltzmann constant, extracted from the fluctuation–dissipation relation $k_{B}=m\hat{\sigma}^{2}/(2\hat{\gamma}T)$ [5, 22], was recovered as $\hat{k}_{B}T/(k_{B}T)_{\mathrm{true}}=0.984$ with bootstrap 95% CI [0.979, 0.989][0.979,\,0.989] from 50 independent simulations (Fig. 4f), a $1.56\%$ error. The drift channel yielded $\hat{\gamma}/\gamma_{\mathrm{true}}=0.971$ with bootstrap 95% CI [0.902, 0.998][0.902,\,0.998] (Fig. 4e). The wider drift-channel CI reflects the lower signal-to-noise ratio of local drift estimation relative to diffusion recovery at these parameters.
@@ 3.3 State-Dependent Diffusion
The pipeline recovered a state-dependent diffusion profile $\hat{\sigma}^{2}(\mu)$ lying on the Poisson line $\sigma^{2}=\mu$ across a two-decade range of mean count levels (Fig. 5b). At a single count rate ($\mu=3000$), the pipeline-recovered diffusion was $\hat{\sigma}/\sqrt{\mu}=1.059$ with bootstrap 95% CI [1.051, 1.062][1.051,\,1.062] from 50 independent pipeline runs, a $5.9\%$ positive bias consistent with the $O(k/N)$ bias of $k$-NN Kramers–Moyal estimators at $k=200$, $N=5000$. Across the variance–mean sweep, this multiplicative bias is approximately uniform and cancels in the log–log slope: the variance–mean exponent was $0.999$ (theoretical $1.000$; error $0.09\%$; $R^{2}=1.0000$). The 250-realisation ensemble achieved 95.8% CI coverage.
@@
The pipeline recovered $\hat{\sigma}/\sigma_{\mathrm{true}}=0.997$ with bootstrap 95% CI [0.995, 1.000][0.995,\,1.000] from 50 independent runs (Fig. 6b), a $0.27\%$ error. For oscillatory systems, the embedding lag $\tau_{\mathrm{embed}}$ is set to the autocorrelation function zero-crossing ($\tau_{\mathrm{ACF}}\approx 155$$160$ steps) rather than unity, consistent with standard practice in nonlinear time series analysis for oversampled oscillatory data, ensuring that the delay-coordinate embedding spans a meaningful portion of the oscillator’s phase space. Across nine frequency values spanning a decade, the recovered $\hat{\sigma}^{2}$ was proportional to $\omega$ with $R^{2}=0.9999$ (Fig. 6c), confirming the linear frequency dependence predicted by the fluctuation–dissipation relation.
@@ 3.4 State-Dependent Drift and Diffusion
The pipeline was applied independently at seven system sizes ($\Omega\in[50,5000]$), with 50 independent simulations per system size. On a log–log plot of recovered $\hat{\sigma}$ versus $\Omega$ (Fig. 7d), each simulation yielded a per-run scaling exponent; the median across 50 runs was $\hat{\alpha}=-0.502$ with bootstrap 95% CI [0.503,0.500][-0.503,\,-0.500] (theoretical $\alpha=-0.500$; error $0.38\%$; $R^{2}=1.0000$). The bootstrap CI contains the theoretical value 0.500-0.500, demonstrating that the multiplicative $k$-NN estimator bias, which is approximately $3\%$ at each individual $\Omega$ value, cancels in the log–log slope (Fig. 7e). The 250-realisation ensemble at $\Omega=500$ achieved 94.0% CI coverage.
@@ 3.5 Relativistic Corrections
For C6, a classical electromagnetic wavepacket was propagated, and the centroid position tracked. The pipeline recovered $\hat{\sigma}/\mathrm{signal}=4.3\times 10^{-7}$ (indistinguishable from zero) and a propagation speed $\hat{c}/c_{\mathrm{true}}=1.0000$ with bootstrap 95% CI [1.0000, 1.0000][1.0000,\,1.0000] from 50 independent runs (Fig. 8a,b).
@@
On a log–log plot of recovered $\hat{\sigma}/\hat{\sigma}_{0}$ versus $\gamma$ (Fig. 8d), the median slope across 50 independent simulations was 0.993-0.993 with bootstrap 95% CI [0.999,0.988][-0.999,-0.988] (theoretical 1.000-1.000; error $0.72\%$; $R^{2}=0.9999$). The per-momentum $\hat{\sigma}/\sigma_{\mathrm{true}}$ ratio was approximately constant at 0.960.960.970.97 across all Lorentz factors, confirming that the $k$-NN estimator bias is multiplicative and cancels in the slope.
@@
The two factors operate at different levels (noise amplitude versus drift), arise from different inputs (special relativity versus Klein–Gordon current conservation), and use different mathematical tools (DDS theorem versus Fokker–Planck self-consistency). Their product $\sigma_{\mathrm{KG}}=\sigma_{0}/\gamma$ is confirmed by the C7 recovery: slope 0.993-0.993 with bootstrap 95% CI [0.999,0.988][-0.999,-0.988], consistent with the predicted exponent 1-1 at the $0.72\%$ level. The independent recovery of the group velocity $v_{g}$ at each momentum mitigates the potential circularity that $\gamma$ depends on $v$: both channels: drift and diffusion: carry relativistic content, recovered independently from the same blind pipeline.